Let X denote the distance (m) that an animal moves from its birth site to the first territorial vacancy it encounters. Suppose that for banner-tailed kangaroo rats, X has an exponential distribution with parameter ? = 0.01327.

(a) What is the probability that the distance is at most 100 m? At most 200 m? Between 100 and 200 m? (Round your answers to four decimal places.)

at most 100 m

at most 200 m

between 100 and 200 m


(b) What is the probability that distance exceeds the mean distance by more than 2 standard deviations? (Round your answer to four decimal places.)


(c) What is the value of the median distance? (Round your answer to two decimal places.)
Answer #1
Concepts and reason

The concept of the probability and the exponential distribution are used to solve this problem.

Integrate the exponential function within the provided limits to determine the required probability.

Fundamentals

An exponential distribution is a continuous distribution.

A continuous random variable X is said to have an exponential distribution with a parameter (λ>0)\left( {\lambda > 0} \right) , if its probability distribution function is given by,

fX(x)={λeλxx00<0{f_X}\left( x \right) = \left\{ \begin{array}{l}\\\lambda {e^{ - \lambda x}}{\rm{ }}x \ge 0\\\\0{\rm{ }} < {\rm{0}}\\\end{array} \right.

Here. λ\lambda is the exponential distribution parameter.

The mean and standard deviation of the exponential distribution is given as,

μ=1λσ=1λ\begin{array}{l}\\\mu = \frac{1}{\lambda }\\\\\sigma = \frac{1}{\lambda }\\\end{array}

Here, μ\mu is the mean of the function and σ\sigma is the standard deviation.

(a.1)

Obtain the exponential distribution parameter from the problem.

λ=0.01327\lambda = 0.01327

Write the probability distribution function.

fX(x)=λeλx{f_X}\left( x \right) = \lambda {e^{ - \lambda x}}

Integrate above function with respect to distance (x)\left( x \right) and take limits from 00 to 100100 .

P(X100)=0100λeλxdx=(1e100λ)=1e100(0.01327)=0.7347\begin{array}{c}\\P\left( {X \le 100} \right) = \int\limits_0^{100} {\lambda {e^{ - \lambda x}}} dx\\\\ = \left( {1 - {e^{ - 100\lambda }}} \right)\\\\ = 1 - {e^{ - 100\left( {0.01327} \right)}}\\\\ = 0.7347\\\end{array}

(a.2)

Write the probability distribution function.

fX(x)=λeλx{f_X}\left( x \right) = \lambda {e^{ - \lambda x}}

Integrate the above function with respect to distance (x)\left( x \right) and take limits from 00 to 200200 .

P(X200)=0200λeλxdx=(1e200λ)=1e200(0.01327)=0.9296\begin{array}{c}\\P\left( {X \le 200} \right) = \int\limits_0^{200} {\lambda {e^{ - \lambda x}}} dx\\\\ = \left( {1 - {e^{ - 200\lambda }}} \right)\\\\ = 1 - {e^{ - 200\left( {0.01327} \right)}}\\\\ = 0.9296\\\end{array}

(a.3)

Write the probability distribution function.

fX(x)=λeλx{f_X}\left( x \right) = \lambda {e^{ - \lambda x}}

Integrate the above function with respect to distance (x)\left( x \right) and take limits 100100 to 200200 .

P(100X200)=100200λeλxdx=(e100λe200λ)=e100(0.01327)e200(0.01327)=0.1949\begin{array}{c}\\P\left( {100 \le X \le 200} \right) = \int\limits_{100}^{200} {\lambda {e^{ - \lambda x}}} dx\\\\ = \left( {{e^{ - 100\lambda }} - {e^{ - 200\lambda }}} \right)\\\\ = {e^{ - 100\left( {0.01327} \right)}} - {e^{ - 200\left( {0.01327} \right)}}\\\\ = 0.1949\\\end{array}

(b)

Write the expression to calculate the mean.

μ=1λ\mu = \frac{1}{\lambda }

Substitute 0.013270.01327 for λ\lambda .

μ=10.01327=75.3579\begin{array}{c}\\\mu = \frac{1}{{0.01327}}\\\\ = 75.3579\\\end{array}

Since the expression of the mean and standard deviation of the exponential distribution are same. Therefore,

σ=μ=75.3549\begin{array}{c}\\\sigma = \mu \\\\ = 75.3549\\\end{array}

Calculate the probability that distance exceeding the mean by more than two standard deviations.

P(Xμ>2σ)=P(X>μ+2σ)=1P(X<μ+2σ)=1FX(μ+2σ)\begin{array}{c}\\P\left( {X - \mu > 2\sigma } \right) = P\left( {X > \mu + 2\sigma } \right)\\\\ = 1 - P\left( {X < \mu + 2\sigma } \right)\\\\ = 1 - {F_X}\left( {\mu + 2\sigma } \right)\\\end{array}

Further solve as,

=1FX(10.01327+20.01327)=1FX(226.07)=e0.01327×226.07=0.04978\begin{array}{c}\\ = 1 - {F_X}\left( {\frac{1}{{0.01327}} + \frac{2}{{0.01327}}} \right)\\\\ = 1 - {F_X}\left( {226.07} \right)\\\\ = {e^{ - 0.01327 \times 226.07}}\\\\ = 0.04978\\\end{array}

(c)

Calculate the value of median distance.

P(XMd)=0.5FX(Md)=0.5eλ×Md=0.5\begin{array}{c}\\P\left( {X \le {M_d}} \right) = 0.5\\\\{F_X}\left( {{M_d}} \right) = 0.5\\\\{e^{ - \lambda \times {M_d}}} = 0.5\\\end{array}

Take log\log on both sides and solve.

λ×Md=ln(0.5)Md=(ln(0.5)λ)\begin{array}{l}\\ - \lambda \times {M_d} = {\rm{ln}}\left( {0.5} \right)\\\\{M_d} = - \left( {\frac{{{\rm{ln}}\left( {0.5} \right)}}{\lambda }} \right)\\\end{array}

Substitute 0.013270.01327 for λ\lambda .

Md=(ln(0.5)0.01327)=52.2341m\begin{array}{c}\\{M_d} = - \left( {\frac{{{\rm{ln}}\left( {0.5} \right)}}{{0.01327}}} \right)\\\\ = 52.2341{\rm{ m}}\\\end{array}

Ans: Part a.1

The probability of the distance at most 100m100{\rm{ m}} is 0.73470.7347 .

Part a.2

The probability of distance at most 200m200{\rm{ m}} is 0.92960.9296 .

Part a.3

The probability of distance lies between 100m100{\rm{ m}} to 200m200{\rm{ m}} is 0.19490.1949 .

Part b

The probability of the distance exceeding the mean by more than two standard deviations is 0.049780.04978 .

Part b

The value of the median distance is 52.2341m52.2341{\rm{ m}} .